# The Interplay of Heterogeneity and Product Detachment in Templated Polymer Copying

### v1

@@ Appendix B Derivation of Coarse-Grained Propensities
Equations 37, 38 and 39 are consistent with the coarse-grained propensities in Equations 8 and 9. Now refer to Figure 4(a) for the case of templated self-assembly without detachment. $\Lambda^{+}_{\textnormal{TSA},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ (Equation 40) and $\Lambda^{-}_{\textnormal{TSA},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ (Equation 41) can again be obtained by a straightforward multiplication of forward and backward rates
$\Lambda^{+}_{\textnormal{TSA},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ $=K^{+}_{\textnormal{bind}}(n_{l-1}n_{l},m_{l-1}m_{l})K^{+}_{\textnormal{pol}}(n_{l-1}n_{l},m_{l-1}m_{l})$
@@
$\Lambda^{-}_{\textnormal{TSA},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ $=K^{-}_{\textnormal{bind}}(n_{l-1}n_{l},m_{l-1}m_{l})K^{-}_{\textnormal{pol}}(n_{l-1}n_{l},m_{l-1}m_{l})$
@@
For $A_{\textnormal{TSA},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$, note that the only two spanning trees for a modified process with polymerization redirected back towards the initial state are the single-step processes of unbinding and polymerization, leading to the form in Equation 42.
$A_{\textnormal{TSA},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ $=K^{-}_{\textnormal{bind}}(n_{l-1}n_{l},m_{l-1}m_{l})+K^{+}_{\textnormal{pol}}(n_{l-1}n_{l},m_{l-1}m_{l})$
@@ Appendix C Derivation of Average Tip Waiting Times
Analogous definitions and abbreviations can be applied to TSA without detachment. The function $\chi_{\textnormal{TSA}}(m_{l+1})=1+m_{l+1}$ now maps each $m_{l+1}$ to $f_{\textnormal{exit}}$, its final state before transition. The flux $J_{\textnormal{TSA}}$ is then calculated as follows (refer to Figure 4(b) for state indices):
$J_{\textnormal{TSA}}=K_{\textnormal{bind}}^{-}(-)p_{ss,\textnormal{TSA}}(1)$
$+\Sigma_{m_{l+1}}K^{+}_{\textnormal{pol}}(m_{l+1})p_{ss}(\chi_{\textnormal{TSA}}(m_{l+1})).$ (47)
Steady state probabilities and normalization variables for TSA without detachment, calculated through standard methods, are given in equations 48, 49 and 50.
$p_{ss,\textnormal{TSA}}(\chi_{\textnormal{TSA}}(m_{l+1}))=\frac{K^{+}_{\textnormal{bind}}(m_{l+1})}{\mathcal{N}_{\textnormal{TSA}}(K_{\textnormal{bind}}^{-}(m_{l+1})+K_{\textnormal{pol}}^{+}(m_{l+1}))},$ (48)
$p_{ss,\textnormal{TSA}}(1)=\frac{K^{-}_{\textnormal{pol}}(-)}{\mathcal{N}_{\textnormal{TSA}}(K_{\textnormal{pol}}^{+}(-)+K_{\textnormal{bind}}^{-}(-))},$ (49)
$\mathcal{N}_{\textnormal{TSA}}$ $=1+\frac{K^{-}_{\textnormal{pol}}(-)}{K_{\textnormal{pol}}^{+}(-)+K_{\textnormal{bind}}^{-}(-)}+\Sigma_{m_{l+1}}\frac{K^{+}_{\textnormal{bind}}(m_{l+1})}{K_{\textnormal{bind}}^{-}(m_{l+1})+K_{\textnormal{pol}}^{+}(m_{l+1})}.$ (50)
@@ Appendix E Parameter Sweep of Relative Error and Total Visitations
Figure 14: Heat map illustrating the resulting relative error change $\langle\ln{\frac{\epsilon}{\epsilon_{I}}}\rangle$ from parameter sweeps for heterogeneity, having extremised over $p_{t}$. We consider heterogeniety on correct (backward propensity discrimination: (a); forward discrimination propensity discrimination: (b)) and incorrect (backward propensity discrimination: (c); forward propensity discrimination: (d)) monomers. In the case of heterogeniety on correct monomer interactions, this relative error change is minimised over $p_{t}$; for heterogeneity on incorrect monomer interactions it is maximised over $p_{t}$. Only the upper triangle $\Delta G_{TT,H}\geq\Delta G_{TT,L}$ is plotted. For heterogeneity on correct monomers, $\min_{p_{t}}\langle\ln{\frac{\epsilon}{\epsilon_{I}}}\rangle=0$ so relative error decreases do not occur, and for heterogeneity on incorrect monomers, $\max_{p_{t}}\langle\ln{\frac{\epsilon}{\epsilon_{I}}}\rangle=0$ so relative error increases do not occur.
Figure 15: Heat map illustrating the resulting log heterogeneous change in state visits $\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}$ from parameter sweeps for heterogeneity, having extremised over $p_{t}$. We consider heterogeniety on correct (backward propensity discrimination: (a); forward discrimination propensity discrimination: (b)) and incorrect (backward propensity discrimination: (c); forward propensity discrimination: (d)) monomers. In the case of heterogeniety on correct monomer interactions, $\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}$ is minmised over $p_{t}$; for heterogeneity on incorrect monomer interactions it is maximised over $p_{t}$. Only the upper triangle $\Delta G_{TT,H}\geq\Delta G_{TT,L}$ is plotted. For heterogeneity on correct monomers,$\min_{p_{t}}\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}=0$ so decreases in state visits do not occur. For heterogeneity on incorrect monomers, $\max_{p_{t}}\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}=0$ for the vast majority of parameters, so increases in state visits do not occur. However, deviations $\max_{p_{t}}\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}>0$ are observed in the small discrimination regime for heterogeneity on incorrect monomer interactions.
@@ Appendix F Investigations on Error Reduction Through Heterogeneity for Separating Copiers
We list our constrained parameter sets in Table 4. For each constrained set, we allow for baseline amounts of forward and backward propensity discrimination $\Delta G_{K}^{*}$ and $\Delta G_{TT}^{*}$ to vary independently. Two values of $\Delta G_{\textnormal{pol}}$ are considered, 00 and $-0.3$ (note that $-0.3$ is weaker driving). For each parameter set, we consider the average error ratio $\mathbb{E}_{p_{t}}[{\frac{\langle\epsilon\rangle}{\langle\epsilon_{I}\rangle}}]$ (here the expectation is taken over a uniform distribution of the parameter $p_{t}$), a measure of how much heterogeniety helps decrease errors averaged over the content of monomer $2$ for Bernoulli templates. Sweeping over different baseline values $G_{K}^{*}$ and $G_{TT}^{*}$, we obtain heat maps of $\mathbb{E}_{p_{t}}[{\frac{\langle\epsilon\rangle}{\langle\epsilon_{I}\rangle}}]$, plotted in figure 16.

### v2

@@ Appendix B Derivation of Coarse-Grained Propensities
Equations 37, 38 and 39 are consistent with the coarse-grained propensities in Equations 8 and 9. Now refer to Figure 4(a) for the case of non-separating templated copolymerization. $\Lambda^{+}_{\textnormal{NTC},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ (Equation 40) and $\Lambda^{-}_{\textnormal{NTC},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ (Equation 41) can again be obtained by a straightforward multiplication of forward and backward rates
$\Lambda^{+}_{\textnormal{NTC},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ $=K^{+}_{\textnormal{bind}}(n_{l-1}n_{l},m_{l-1}m_{l})K^{+}_{\textnormal{pol}}(n_{l-1}n_{l},m_{l-1}m_{l})$
@@
$\Lambda^{-}_{\textnormal{NTC},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ $=K^{-}_{\textnormal{bind}}(n_{l-1}n_{l},m_{l-1}m_{l})K^{-}_{\textnormal{pol}}(n_{l-1}n_{l},m_{l-1}m_{l})$
@@
For $A_{\textnormal{NTC},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$, note that the only two spanning trees for a modified process with polymerization redirected back towards the initial state are the single-step processes of unbinding and polymerization, leading to the form in Equation 42.
$A_{\textnormal{NTC},\mathbf{x}}(n_{l-1}n_{l},m_{l-1}m_{l})$ $=K^{-}_{\textnormal{bind}}(n_{l-1}n_{l},m_{l-1}m_{l})+K^{+}_{\textnormal{pol}}(n_{l-1}n_{l},m_{l-1}m_{l})$
@@ Appendix C Derivation of Average Tip Waiting Times
Analogous definitions and abbreviations can be applied to NTC. The function $\chi_{\textnormal{NTC}}(m_{l+1})=1+m_{l+1}$ now maps each $m_{l+1}$ to $f_{\textnormal{exit}}$, its final state before transition. The flux $J_{\textnormal{NTC}}$ is then calculated as follows (refer to Figure 4(b) for state indices):
$J_{\textnormal{NTC}}=K_{\textnormal{bind}}^{-}(-)p_{ss,\textnormal{NTC}}(1)$
$+\Sigma_{m_{l+1}}K^{+}_{\textnormal{pol}}(m_{l+1})p_{ss}(\chi_{\textnormal{NTC}}(m_{l+1})).$ (47)
Steady state probabilities and normalization variables for NTC without detachment, calculated through standard methods, are given in equations 48, 49 and 50.
$p_{ss,\textnormal{NTC}}(\chi_{\textnormal{NTC}}(m_{l+1}))=\frac{K^{+}_{\textnormal{bind}}(m_{l+1})}{\mathcal{N}_{\textnormal{NTC}}(K_{\textnormal{bind}}^{-}(m_{l+1})+K_{\textnormal{pol}}^{+}(m_{l+1}))},$ (48)
$p_{ss,\textnormal{NTC}}(1)=\frac{K^{-}_{\textnormal{pol}}(-)}{\mathcal{N}_{\textnormal{NTC}}(K_{\textnormal{pol}}^{+}(-)+K_{\textnormal{bind}}^{-}(-))},$ (49)
$\mathcal{N}_{\textnormal{NTC}}$ $=1+\frac{K^{-}_{\textnormal{pol}}(-)}{K_{\textnormal{pol}}^{+}(-)+K_{\textnormal{bind}}^{-}(-)}+\Sigma_{m_{l+1}}\frac{K^{+}_{\textnormal{bind}}(m_{l+1})}{K_{\textnormal{bind}}^{-}(m_{l+1})+K_{\textnormal{pol}}^{+}(m_{l+1})}.$ (50)
@@ Appendix E Parameter Sweep of Relative Error and Total Visitations
Figure 14: Heat map illustrating the resulting relative error change $\langle\ln{\frac{\epsilon}{\epsilon_{I}}}\rangle$ from parameter sweeps for heterogeneity, having extremised over $p_{t}$. We consider heterogeneity on correct (backward propensity discrimination: (a); forward discrimination propensity discrimination: (b)) and incorrect (backward propensity discrimination: (c); forward propensity discrimination: (d)) monomers. In the case of heterogeneity on correct monomer interactions, this relative error change is minimised over $p_{t}$; for heterogeneity on incorrect monomer interactions it is maximised over $p_{t}$. Only the upper triangle $\Delta G_{TT,H}\geq\Delta G_{TT,L}$ is plotted. For heterogeneity on correct monomers, $\min_{p_{t}}\langle\ln{\frac{\epsilon}{\epsilon_{I}}}\rangle=0$ so relative error decreases do not occur, and for heterogeneity on incorrect monomers, $\max_{p_{t}}\langle\ln{\frac{\epsilon}{\epsilon_{I}}}\rangle=0$ so relative error increases do not occur.
Figure 15: Heat map illustrating the resulting log heterogeneous change in state visits $\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}$ from parameter sweeps for heterogeneity, having extremised over $p_{t}$. We consider heterogeneity on correct (backward propensity discrimination: (a); forward discrimination propensity discrimination: (b)) and incorrect (backward propensity discrimination: (c); forward propensity discrimination: (d)) monomers. In the case of heterogeneity on correct monomer interactions, $\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}$ is minmised over $p_{t}$; for heterogeneity on incorrect monomer interactions it is maximised over $p_{t}$. Only the upper triangle $\Delta G_{TT,H}\geq\Delta G_{TT,L}$ is plotted. For heterogeneity on correct monomers,$\min_{p_{t}}\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}=0$ so decreases in state visits do not occur. For heterogeneity on incorrect monomers, $\max_{p_{t}}\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}=0$ for the vast majority of parameters, so increases in state visits do not occur. However, deviations $\max_{p_{t}}\ln{\frac{\langle V\rangle}{\langle V_{I}\rangle}}>0$ are observed in the small discrimination regime for heterogeneity on incorrect monomer interactions.
@@ Appendix F Investigations on Error Reduction Through Heterogeneity for Separating Copiers
We list our constrained parameter sets in Table 4. For each constrained set, we allow for baseline amounts of forward and backward propensity discrimination $\Delta G_{K}^{*}$ and $\Delta G_{TT}^{*}$ to vary independently. Two values of $\Delta G_{\textnormal{pol}}$ are considered, 00 and $-0.3$ (note that $-0.3$ is weaker driving). For each parameter set, we consider the average error ratio $\mathbb{E}_{p_{t}}[{\frac{\langle\epsilon\rangle}{\langle\epsilon_{I}\rangle}}]$ (here the expectation is taken over a uniform distribution of the parameter $p_{t}$), a measure of how much heterogeneity helps decrease errors averaged over the content of monomer $2$ for Bernoulli templates. Sweeping over different baseline values $G_{K}^{*}$ and $G_{TT}^{*}$, we obtain heat maps of $\mathbb{E}_{p_{t}}[{\frac{\langle\epsilon\rangle}{\langle\epsilon_{I}\rangle}}]$, plotted in figure 16.